zarkam21:

Compute the following linear combinations. a. 4(-2, 3) - 2(6, -7) b. (79.4, 58.1) + 3(-0.8, 6.3)

Vocaloid:

let's take the x-coordinates first a) 4(-2) - 2(6) = ?

zarkam21:

-20

Vocaloid:

awesome, now the y-coordinates 4(3) - 2(-7) = ?

zarkam21:

-2

zarkam21:

26*

Vocaloid:

awesome so putting the x and y together we get (-20,26) as the solution for a) would you like to try b) using the same line of reasoning? I can help you out if you get stuck

zarkam21:

yeah sure i will do it

zarkam21:

79.4+3(-0.8)=77 58.1+3(6.3)=77 (77,77)

Vocaloid:

perfect